OpenStudy (anonymous):

Charles has read 180 pages of a book. This is 3/4 of the total number of pages in the book. If p represents the number of pages in the book, what is p?

OpenStudy (anonymous):

(3/4)p = 180 p = 180/(3/4) = 240 pages

OpenStudy (llort):

180*(4/3)=x

OpenStudy (anonymous):

240 pageslk\